# HEART: Achieving Timely Multi-Model Training for Vehicle-Edge-Cloud-Integrated Hierarchical Federated Learning

### v1

@@ 13            Dynamically adjust $\pi(\tau)$ based on (20)
23                              $\overline{x}_{m,n;[g]}^{global,\left\langle j\right\rangle}\xleftarrow[]{}\overline{x}_{m,n;[g]}^{p,\left\langle j\right\rangle}(\tau)$,$f_{m,n;[g]}^{global}\xleftarrow[]{}f_{m,n;[g]}^{p}(\tau)$
@@ V Numerical Evaluations
The VEC-HFL architecture considered in our simulations contains 1 CS, 5 ESs, and 25 to 50 vehicles depending on the specific simulation setup, where the CS with the coverage of 5 km is the radius in located at the center of the area. We consider an area with 4 roads and 4 intersections where an ES with the coverage of 1 km is located at each intersection. Each vehicle moves at a constant speed in the road, and at each intersection it takes a turn at a direction chosen randomly. To capture diverse training tasks, we consider 4 different ML tasks: CIFAR-10 (trained on VGG16 neural network [24]), MNIST (trained on a convolutional neural network with 4 layers [5]), Driver Yawning (trained on ResNet-18 neural network [25]) and 20 Newsgroups (trained on LSTM neural network [26]). In particular, CIFAR-10 and MNIST are colored and grayscale image datasets for ten-category image classification, respectively[27, 5], Driver Yawning represents a dataset for yawning detection during driving[28], while 20 Newsgroups is a text dataset in natural language processing, with data spanning 20 distinct topics. The data rates of vehicle-to-edge and edge-to-vehicle communications are set according to the channel modeling in [17]. Other key parameters used in the simulations are listed in Table I.
$\bullet$ Two-Stage Particle Swarm Optimization (TSPSO): Task scheduling and training rank are determined by traditional PSO under time constraints and task training balance [18]. The time complexity of the method is $\mathcal{O}(2N\times J\times p^{*}\times\tau^{*})$.
$\bullet$ Two-Stage Genetic Algorithm (TSGA): Task scheduling and training rank and are determined by traditional GA under time constraints and task training balance[19]. The time complexity of the method is $\mathcal{O}(2N\times J\times p^{*}\times\tau^{*})$.
$\bullet$ Two-Stage Greedy Algorithm (TSGD): On the premise of ensuring that the task training time does not exceed the vehicles’ dwell times, the task scheduling score is calculated for each vehicle based on the task weight coefficient $\rho_{j}$ and the number of times it is allocated. The task with a higher score is selected and assigned to the vehicle, and then the task is removed. The above operation continues until the task scheduling is completed. Afterwards, Alg. 3 is employed to further refine the solution. The time complexity of the method is $\mathcal{O}(2N\times J\times K^{\langle j\rangle})$; $K^{\langle j\rangle}$ is less than $p^{*}$ and $\tau^{*}$.
TABLE I: Parameter settings[29, 30]
@@
In Fig. 4, we study the model accuracy performance of various methods. We set the desired accuracy for the 4 tasks at 0.75, 0.9, 0.9, and 0.75, respectively, which are shown through the dotted horizontal lines in subfigures of Fig. 4. These desired accuracies are further used in the subsequent results. Besides, in the legends of the plots, we use ComplT to represent the wall-clock time at the instant of finishing of each each task (e.g., task CIFAR in plot Fig. 4(a)). Also, we highlight the time in which the last task gets completed in each method via the red color in the legends (e.g., as in plot Fig. 4(b), TSSO finishes the processing of all tasks through the execution of MNIST and the finish time is 215.8s). As shown in Fig. 4, our method completes the MNIST task training at 49.8s, the Driver Yawning task training at 74.7s, the CIFAR task training at 99.3s, and the time required to complete four task training is 110.2s, which is when the execution of 20 Newsgroup task completes. An interesting observation from Fig. 4 is that our method does not necessarily lead to the fastest convergence of all the tasks when the convergence speed is measured through the accuracy vs the global iteration. This is mainly due to the fact that our method balances the resource allocation across all the tasks, leading to a uniform convergence speed across all the tasks. Nevertheless, focusing on the values highlighted by red, our method leads to the fastest finish of execution of all tasks when the convergence speed is measured though the wall-clock time (our method finishes all the task in 110.2s, which is notable lower than 191.9s for TSPSO, 154.1s for TSGA, 215.8s for TSSO, and 170.7s for TSGD). This unveils the power of balanced task execution in reducing the overall task completion through similar convergence speeds across different tasks. The rapid convergence of a specific task, as illustrated in Fig. 4(a) with examples like TSPSO and TSGD, is often due to prioritizing that task at the expense of others. In particular, assigning more vehicles to a particular task can indeed improve its convergence rate; however, this can lead to an imbalance across tasks, ultimately compromising equitable performance distribution. For example, in the left subplot of Fig. 3, we can see that the TSGD performs poorly on Driver Yawning, failing to achieve the expected accuracy (Fig. 4(c)). In a summary, our proposed HEART achieves the best convergence speed in terms of wall-clock time across all the baselines (Fig. 4) while keeping a balanced scheduling (Fig. 3).
The performance in terms of time efficiency captured via the overall time overhead (the time for all task models to reach the aforementioned desired accuracies) is shown in Fig. 5, upon testing various numbers of vehicles, where our method HEART, outperforms other methods. For example, we can achieve an average reduction of overall time overhead by $17.2\%$, $24.9\%$, $24.4\%$ and $30.6\%$ as compared to TSGA, TSPSO, TSSO and TSGD in Fig. 5(a) with 25 vehicles; and a reduction in overall time overhead by $18.3\%$, $24.7\%$, $34.59\%$ and $36.76\%$ in Fig. 5(b) when considering 50 vehicles. Interestingly, Fig. 5 shows that time cost is not always affected by the increase in data volume per vehicle. This is because the increase of data volume increases local training time, but also reduces the number of global iterations required to hit the target accuracy.

### v2

@@ 13    Dynamically adjust $\pi(\tau)$ based on (20)
23             $\overline{x}_{m,n;[g]}^{global,\left\langle j\right\rangle}\xleftarrow[]{}\overline{x}_{m,n;[g]}^{p,\left\langle j\right\rangle}(\tau)$$f_{m,n;[g]}^{global}\xleftarrow[]{}f_{m,n;[g]}^{p}(\tau)$
@@ V Numerical Evaluations
The VEC-HFL architecture considered in our simulations contains 1 CS, 5 ESs, and 25 to 50 vehicles depending on the specific simulation setup, where the CS with the coverage of 5 km is the radius in located at the center of the area. We consider an area with 4 roads and 4 intersections where an ES with the coverage of 1 km is located at each intersection. Each vehicle moves at a constant speed in the road, and at each intersection it takes a turn at a direction chosen randomly. To capture diverse training tasks, we consider 4 different ML tasks: CIFAR-10 (trained on VGG16 neural network [VGG16]), MNIST (trained on a convolutional neural network with 4 layers [HiFlash]), Driver Yawning (trained on ResNet-18 neural network [Resnet18]) and 20 Newsgroups (trained on LSTM neural network [LSTM]). In particular, CIFAR-10 and MNIST are colored and grayscale image datasets for ten-category image classification, respectively[Privacy, HiFlash], Driver Yawning represents a dataset for yawning detection during driving[JHPFA], while 20 Newsgroups is a text dataset in natural language processing, with data spanning 20 distinct topics. The data rates of vehicle-to-edge and edge-to-vehicle communications are set according to the channel modeling in [Platoon]. Other key parameters used in the simulations are listed in Table II.
TABLE II: Parameter settings[luo2020hfel, KHEF]
@@
$\bullet$ Two-Stage Particle Swarm Optimization (TSPSO): Task scheduling and training rank are determined by traditional PSO under time constraints and task training balance [DynamicSR]. The time complexity of the method is $\mathcal{O}(2N\times J\times p^{*}\times\tau^{*})$.
$\bullet$ Two-Stage Genetic Algorithm (TSGA): Task scheduling and training rank and are determined by traditional GA under time constraints and task training balance[Swarm, OffloadedT]. The time complexity of the method is $\mathcal{O}(2N\times J\times p^{*}\times\tau^{*})$.
$\bullet$ Two-Stage Greedy Algorithm (TSGD): On the premise of ensuring that the task training time does not exceed the vehicles’ dwell times, the task scheduling score is calculated for each vehicle based on the task weight coefficient $\rho_{j}$ and the number of times it is allocated. The task with a higher score is selected and assigned to the vehicle, and then the task is removed. The above operation continues until the task scheduling is completed. Afterwards, Alg. 3 is employed to further refine the solution [tang2023fedml]. The time complexity of the method is $\mathcal{O}(2N\times J\times K^{\langle j\rangle})$; $K^{\langle j\rangle}$ is less than $p^{*}$ and $\tau^{*}$.
@@
In Fig. 4, we study the model accuracy performance of various methods. We set the desired accuracy for the 4 tasks at 0.75, 0.9, 0.9, and 0.75, respectively, which are shown through the dotted horizontal lines in subfigures of Fig. 4. These desired accuracies are further used in the subsequent results. Besides, in the legends of the plots, we use ComplT to represent the wall-clock time at the instant of finishing of each each task (e.g., task CIFAR in plot Fig. 4(a)). Also, we highlight the time in which the last task gets completed in each method via the red color in the legends (e.g., as in plot Fig. 4(b). TSSO finishes the processing of all tasks through the execution of MNIST and the finish time is 215.8s). As shown in Fig. 4, our method completes the MNIST task training at 49.8s, the Driver Yawning task training at 74.7s, the CIFAR task training at 99.3s, and the time required to complete four task training is 110.2s, which is when the execution of 20 Newsgroup task completes. An interesting observation from Fig. 4 is that our method does not necessarily lead to the fastest convergence of all the tasks when the convergence speed is measured through the accuracy vs the global iteration. This is mainly due to the fact that our method balances the resource allocation across all the tasks, leading to a uniform convergence speed across all the tasks.
@@
Nevertheless, focusing on the values highlighted by red, our method leads to the fastest finish of execution of all tasks when the convergence speed is measured though the wall-clock time (our method finishes all the task in 110.2s, which is notable lower than 191.9s for TSPSO, 154.1s for TSGA, 215.8s for TSSO, and 170.7s for TSGD). This unveils the power of balanced task execution in reducing the overall task completion through similar convergence speeds across different tasks. The rapid convergence of a specific task, as illustrated in Fig. 4(a) with examples like TSPSO and TSGD, is often due to prioritizing that task at the expense of others. In particular, assigning more vehicles to a particular task can indeed improve its convergence rate; however, this can lead to an imbalance across tasks, ultimately compromising equitable performance distribution. For example, in the left subplot of Fig. 3, we can see that the TSGD performs poorly on Driver Yawning, failing to achieve the expected accuracy (Fig. 4(c)). In a summary, our proposed HEART achieves the best convergence speed in terms of wall-clock time across all the baselines (Fig. 4) while keeping a balanced scheduling (Fig. 3).
The performance in terms of time efficiency captured via the overall time overhead (the time for all task models to reach the aforementioned desired accuracies) is shown in Fig. 5, upon testing various numbers of vehicles, where our method HEART, outperforms other methods. For example, we can achieve an average reduction of overall time overhead by 17.2%, 24.9%, 24.4% and 30.6% as compared to TSGA, TSPSO, TSSO and TSGD in Fig. 5(a) with 25 vehicles; and a reduction in overall time overhead by 18.3%, 24.7%, 34.59% and 36.76% in Fig. 5(b) when considering 50 vehicles. Interestingly, Fig. 5 shows that time cost is not always affected by the increase in data volume per vehicle. This is because the increase of data volume increases local training time, but also reduces the number of global iterations required to hit the target accuracy.